State True or False : If for a given current,resistance is changed from R to R/n,powerconsumed changes from P to nP.
1
Expert's answer
2012-07-05T10:10:42-0400
P = V · I = (I · R) · I = I^2 · R
P0 = I^2 · R
P1 = I^2 · R /
n

Answer: false, power consumed changes from P to P/n.